OpenStudy (anonymous):

simplify.. having a hard time with radicals :( 1/2√1-x^2•1/√1-√1-x^2

OpenStudy (kamille):

can you write this using latex?

OpenStudy (anonymous):

-2x^2-1/2

OpenStudy (anonymous):

or -2(x+(sqrt1/2))(1-(sqrt(1/2))

OpenStudy (jmark):

Given 1/2√1-x^2•1/√1-√1-x^2 simplification of this =-2x²-1/2

OpenStudy (anonymous):

*R:R means under radical ! 1(R)=1 just it !:) First answer to :- x^2.1/1(R)=-x^2.1 Then : 1/2√1 = 1/2 And then : -√1-x^2=-1-x^2 So we have : 1/2√1-x^2•1/√1-√1-x^2 = 1/2 - 1 - x^2 - x^2 = -1+1/2 = -1/2-2x^2 Just this :) @koleendiaz : If u have any problem send a massage :)